40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Dec 7 13:11:57 2014, manasa, Summary, General, Frequency Offset Locking - To Do List FOLtodolist.pdf
    Reply  Mon Dec 8 15:54:39 2014, manasa, Summary, General, Dec 8 - Check Frequency Counter module 
       Reply  Mon Dec 8 20:53:51 2014, diego, Summary, General, Dec 8 - Check Frequency Counter module 
          Reply  Tue Dec 9 16:06:46 2014, diego, Summary, General, Dec 8 - Check Frequency Counter module 
    Reply  Tue Dec 9 00:30:27 2014, manasa, Summary, General, Dec 9 - Elaborate to do list FOL_FiberChassis.pdf
       Reply  Tue Dec 9 16:07:16 2014, manasa, Summary, General, Dec 9 - FC module and fiber chassis FCmodule.png
    Reply  Wed Dec 10 16:12:29 2014, manasa, Summary, General, Dec 10 - PSL table FiberMod_front.jpgFOL_fiber.jpg
       Reply  Thu Dec 11 12:50:12 2014, manasa, Summary, General, PSL table optical layout 
          Reply  Thu Dec 11 23:34:06 2014, manasa, Summary, General, PSL table optical layout 
             Reply  Fri Dec 12 02:30:25 2014, Jenne, Summary, General, PSL table optical layout 
    Reply  Fri Dec 12 14:38:39 2014, manasa, Summary, General, Frequency Offset Locking - To Do List (Revised) FOLtodolist.pdf
       Reply  Fri Dec 12 15:19:04 2014, manasa, Summary, General, Dec 12 - PSL table PSLtoFiber.pngPSLfiberChassis.pngPSLfiberChassis.png
       Reply  Wed Dec 17 11:57:56 2014, manasa, Summary, General, Y arm optical layout 58.png
Message ID: 10808     Entry time: Wed Dec 17 11:57:56 2014     In reply to: 10791
Author: manasa 
Type: Summary 
Category: General 
Subject: Y arm optical layout 

I was working around the PSL table and Y endtable today.

I modified the Y arm optical layout that couples the 1064nm light leaking from the SHG crystal into the fiber for frequency offset locking.

The ND filter that was used to attenuate the power coupled into the fiber has been replaced with a beam sampler (Thor labs BSF-10C). The reflected power after this optic is ~1.3mW and the trasmitted power has been dumped to a razor blade beam dump (~210mW).

Since we have a spare fiber running from the Y end  to the PSL table, I installed an FC/APC fiber connector on the PSL table to connect them and monitored the output power at the Y end itself. After setting up, I have ~620uW of Y arm light on the PSL table (~48% coupling).

During the course of the alignment, I lowered the power of the Y end NPRO and disengaged the ETMY oplev. These were reset after I closed the end table.

Attached is the out of loop noise measurement of the Y arm ALS error signal before (ref plots) and after.

 

Attachment 1: 58.png  81 kB  | Hide | Hide all
58.png
ELOG V3.1.3-